Aluno  Matrícula Nota
Alexandre Vervloet  99900234 9,4
Arthur Antonio Brabosa SOARES JR 99900278 5,5
Fabrício Vazole  99900238 6,5
Francisco de Assis Favarato Cutini  99900185 5,5
Hilário Seibel Junior  99900247 8,9
Jaci Alves de souza JR  99900269 7,3
Jociel Cavalcante Andrade  99900269 8,9
Kellen Suamy Salla Barcelos  99900245 6,9
Luciana Lee  99901466 7,0
Márcio da Rós Gomes  99900177 9,6
Mariella Berger 9990282 7,6
Michelle Siqueira Stange  99900237 5,5
Otávio Martins Carneiro  99900173 7,3
Patrick Barbosa da Silva  99900266 7,5
Suzana Dornelas Bassani  99900273 5,5

1. Estruturas de dados

Aluno  Matrícula Nota
Daniel Guimarães Santos  9904329  7,1
Gabriela Trazi Perim  9903165  9,1
Letícia Rossetti Margoto  99903128 9,2
Felipe Gomes Carvalho  99903095 8,6
Rodrigo Alves Sarmento  99903066 6,5
Silvano Nogueira Buback  99903088 6,5
Suellen Marconsini Loureiro 99903120 8,6